OpenStudy (anonymous):

Sam is observing the velocity of a car at different times. After two hours, the velocity of the car is 50 km/h. After six hours, the velocity of the car is 54 km/h. Part A: Write an equation in two variables in the standard form that can be used to describe the velocity of the car at different times. Show your work and define the variables used. (5 points) Part B: How can you graph the equation obtained in Part A for the first seven hours? (5 points)

OpenStudy (anonymous):

Please help me

OpenStudy (s.gracex):

so i think this is a linear equation. so the formula is y=mx+b so, m = rise/run =(y2-y1)/(x2-x1) the x's and y's are the points given: "After three hours, the velocity of the car is 53 km/h. After six hours, the velocity of the car is 62 km/h" (x1,y1) = (2,50) (x2,y2) = (6,54) values back into the equation m = (54-50)/(6-2) m = 4/4 m = 0 now we use a point-slope form to find the the standard form y-y1 = m(x-x1) where x1 and y1 are any set of point given y-50 = 2(x-2) y-50 = 2x - 4 y = 2x - 4 + 50 y = 2x + 46 y is the velocity of the car, x is the time

OpenStudy (anonymous):

what about part b

OpenStudy (s.gracex):

Sorry about that. Ok so what you'll do is draw a graph and your x coordinate will be 0 and your y will be 48, so then you will just mark a strait line all the way across at 0,48

OpenStudy (anonymous):

OK then after that what do i do

OpenStudy (anonymous):

@s.gracex please help me

OpenStudy (anonymous):

@s.gracex please help me i really need your help

OpenStudy (anonymous):

@s.gracex I really need your help, please help me